### v1

@@ Abstract
Singularly perturbed dynamical systems, commonly known as fast-slow systems, play a crucial role in various applications such as plasma physics. They are closely related to reduced order modeling, closures, and structure-preserving numerical algorithms for multiscale modeling. A powerful and well-known tool to address these systems is the Fenichel normal form, which significantly simplifies fast dynamics near slow manifolds through a transformation. However, the Fenichel normal form is difficult to realize in conventional numerical algorithms. In this work, we explore an alternative way of realizing it through structure-preserving machine learning. Specifically, a fast-slow neural network (FSNN) is proposed for learning data-driven models of singularly perturbed dynamical systems with dissipative fast timescale dynamics. Our method enforces the existence of a trainable, attracting invariant slow manifold as a hard constraint. Closed-form representation of the slow manifold enables efficient integration on the slow time scale and significantly improves prediction accuracy beyond the training data. We demonstrate the FSNN on several examples that exhibit multiple timescales, including the Grad moment system from hydrodynamics, two-scale Lorentz96 equations for modeling atmospheric dynamics, and Abraham-Lorentz dynamics modeling radiation reaction of electrons in a magnetic field.
@@
Singularly perturbed dynamical systems, slow manifold, model discovery, closures
@@ 1 Introduction
In this article, a new architecture called the Fast-Slow Neural Network (FSNN) is developed for learning data-driven models of fast-slow dynamical systems with dissipative fast timescale dynamics. Our method enforces the existence of a trainable, attracting invariant slow manifold [10] as a hard constraint. The constraint is imposed by way of a novel neural network architecture with structured weight matrices. By leveraging the neural ODE framework from [11], our technique produces deterministic evolution equations that approximately reproduce trajectory training data. When the training data comes from an unknown dynamical system, the method achieves data-driven full-order model discovery with stability guarantees due to provable existence of the slow manifold. In short, our architecture strongly contains an attractor of singularly perturbed dynamical systems by design.
@@
Nonlinear dimensionality reduction inspired by inertial manifold theory has a rich history (see [36] and the references therein) originating in the 1980s and 1990s [22, 24, 23, 33, 56, 57, 31, 28]. Data-driven approaches for discovering inertial manifolds in dynamical systems and using separations of scales to evolve a lower-dimensional reduced order model first materialized in the 90s [34, 59, 37, 38, 55, 53, 54] and recently reemerged in popularity due to boom in machine-learning-based modeling [4, 12, 48, 3, 39, 42, 61, 43, 60, 16, 44, 41]. The work in this manuscript tackles a special class of systems with attracting invariant sets, namely fast-slow systems with dissipative fast timescale dynamics [10]. Unlike previous data-driven approaches, the FSNN is a full-order model which, through the use of neural network architectures with special properties, strictly enforces the existence of an attracting slow manifold. This method may be viewed as a major step along a pathway in developing data-driven model discovery techniques with embedded invariant manifolds that do not use explicit timescale separation as a crutch.
@@
The design of the FSNN architecture is motivated by the the first few theorems discovered by Fenichel [35]. The components of the architecture are constructed using structured weight matrices. The coordinate transformation into Fenichel normal form is based on an invertible coupling flow network [58] in which we introduced the additional structure-preserving property of bi-Lipschitz control through the use of bi-Lipschitz affine transformation (bLAT) layers. This network and its inverse can be evaluated using closed-form expressions with identical computational complexity. We note that “vanilla” neural networks are generally non-invertible [25, 52, 14]. We also introduce the Schur form network, which is a novel architecture based on the Schur decomposition for controlling the locations of eigenvalues in the complex plane, to parameterize the linear term in the dynamics. Additionally, a low-rank bilinear form network is included to increase the approximation power in the fast variable space. Our approach is then integrated using adaptive time stepping that treats the linear term implicitly, further leveraging back-substitution on the Schur form. Some of the structures introduced are related to the idea of trivializing manifold optimization through parameterizations [40]. Structure-preserving flow maps through machine learning are an active area of research, and include such as the symplectic neural networks for Hamiltonian dynamics [32, 9]. Previous studies on flow map learning for multi-scale dynamical systems include [46] and [20]. In particular, [20] simultaneously preserves symplecticity and adiabatic invariance in a multi-scale setting.
Before going into a detailed presentation of our new method, it is worth highlighting an important shortcoming it suffers from, and how that shortcoming might be mitigated in practice. The method assumes that the dimension of the attracting invariant manifold is known in advance. While this is not a limitation in the surrogate modeling context, where the (expensive to simulate) dynamical system is known in detail, it does present a moderate challenge in the model discovery context. As such, when performing data-driven model discovery using our method, it will be necessary to first determine the dimension of the slow manifold, either through physical reasoning, or through application of auxiliary data mining techniques such as diffusion maps [13], autoencoders [61], or other techniques from reduced order modeling [36].
@@ 2 Fast-Slow Dynamics and the Fenichel Normal Form
In this paper, we develop a novel machine learning (ML)-inspired architecture, the Fast-Slow Neural Network (FSNN). FSNNs are a parameterization for a class of dynamical systems exhibiting multiscale behavior in time that, to leading order, are dissipative on the fast time scale. Fast-slow systems, which exhibit both short-term and long-term behavior, are relevant to many physical systems observed in nature, including the climate and plasma systems. In this section, we motivate the FSNN architecture with a review the theory of fast-slow systems.
@@
A fast-slow system is an ODE of the form [10]
$\epsilon\frac{d}{d\tau}y$ $=f(x,y,\epsilon),$ (1a)
$\frac{d}{d\tau}x$ $=g(x,y,\epsilon),$ (1b)
@@
$\epsilon\frac{d}{d\tau}z$ $=U(z,\epsilon),$ (2)
@@
This lemma is proved following Theorem 4 in [35]. The purpose of the FSNN is to be a universal approximator for normally stable fast-slow systems. The network is based on the following well known result due to Fenichel.
@@
$\frac{d}{dt}\tilde{y}$ $=A(\tilde{x},\tilde{y},\epsilon)\tilde{y},$ (4a)
$\frac{d}{dt}\tilde{x}$ $=\epsilon{\tilde{g}}(\tilde{x},\tilde{y},\epsilon),$ (4b)
@@
The above lemma is derived in Chapter 3.2 of [35]. The lemma introduces (2), which is often called the Fenichel normal form due to the pioneer work of Neil Fenichel [21]. Note that the Fenichel normal form given in [35] is more general, including fixed points that are either stable or unstable in the linearization of the fast variable. In this work, we only consider fast variables that are long-time stable. Moreover, due to the existence of the coordinate transformation, we consider the general form of the singularly perturbed systems (2). However, the advantage of the split form (2) is that the slow manifold dimension is known.
@@
$\left[\begin{array}[]{c}y\\ x\end{array}\right]$ $=h(z;\epsilon),$ (7)
@@
$\frac{d}{dt}y$ $=T(x)y+B(x,y,\epsilon)(y,y)+\epsilon C(x,y,\epsilon)y,$ (8a)
$\frac{d}{dt}x$ $=\epsilon g(x,y,\epsilon).$ (8b)
$T(x)\in\mathbb{R}^{N_{y}\times N_{y}}$ is a block upper triangular matrix consisting of either $\mathbb{R}^{2\times 2}$ or $\mathbb{R}$ blocks on the diagonal and the eigenvalues of these diagonal blocks are strictly to the left of the imaginary axis for all $x$. $B(x,y,\epsilon):\mathbb{R}^{N_{y}}\times\mathbb{R}^{N}_{y}\rightarrow\mathbb{R}^{N}_{y}$ is a bilinear map, $C(x,y,\epsilon)\in\mathbb{R}^{N_{y}\times N_{y}}$ is a matrix, and $g(x,y,\epsilon)\in\mathbb{R}^{N_{x}}$. $T$, $B$, $C,$ and $g$ are smooth functions of their arguments.
In the following discussion, we will prove Theorem 2 by refining the Fenichel normal form in (2) and applying the matrix Schur decomposition. First, we modify (2) by performing a Taylor expansion of the term $A(\tilde{x},\tilde{y},\epsilon)\tilde{y}$ about $y=0$ and $\epsilon=0$. With sufficient smoothness on $A$, we have
$A(\tilde{x},\tilde{y},\epsilon)\tilde{y}=A(\tilde{x},0,0)\tilde{y}+B(\tilde{x},\tilde{y},\epsilon)(\tilde{y},\tilde{y})+\epsilon C(\tilde{x},\tilde{y},\epsilon)\tilde{y},$ (9)
@@
The Schur form of a matrix, $A\in\mathbb{R}^{M\times M}$, is a matrix $T\in\mathbb{R}^{M\times M}$ that satisfies $A=QTQ^{T}$, where $Q\in\mathbb{R}^{M\times M}$ is orthogonal ($Q^{T}Q=I$) and $T$ is block upper triangular where the diagonal blocks consist of $\mathbb{R}$ and $\mathbb{R}^{2\times 2}$ blocks.
[…]
The Schur decomposition of a matrix $A\in\mathbb{R}^{M\times M}$, is given by
$A=QTQ^{\rm T},$ (10)
where $Q,T\in\mathbb{R}^{M\times M}$, $Q$ is an orthogonal matrix ($Q^{\rm T}Q=I$), and $T$ is the Schur form of $A$. The eigenvalues of $A$ are given by the eigenvalues of the diagonal blocks of $T$.
The existence of the Schur decomposition is proven in many references, including [30, 2], though for any given matrix, the decomposition is not unique. Let the Schur decomposition of $A(\tilde{x},0,0)$ be given by
$A(\tilde{x},0,0)=Q(\tilde{x})T(\tilde{x})Q(\tilde{x})^{\rm T},$ (11)
@@
where $\tilde{B}(\tilde{x},\hat{y},\epsilon)$ is a bilinear form operator and $\tilde{C}(\tilde{x},\hat{y},\epsilon)$ is a matrix. This establishes an equivalence to (2).
@@
$\frac{d}{d\tau}x=g(x,0,\epsilon),\qquad(x,y)\in M,$ (12)
@@
$d\left((x(t),y(t)),M\right)\leq\kappa\exp(\alpha t),$ (13)
@@ 3 Fast-Slow Neural Networks (FSNNs)
The fast slow neural network (FSNN) is a transformation of an initial condition, $z_{0}$, and parameter, $\epsilon>0$, to a time series $\{z_{n}\}_{n=0}^{N}$ corresponding to times $t_{0},\dots,t_{N}$, where $z_{n}\in\mathbb{R}^{N}$ for $n=0,\dots,N$. The initial condition is transformed into fast-slow coordinates, $(y_{0},x_{0})$, via (7) where $h$ is represented using an invertible coupling flow network. A numerical integration is used to approximate the solution to the fast-slow system in (2) to obtain the time series $\{(y_{n},x_{n})\}_{n=0}^{N_{t}}$ for corresponding times $t_{0},\dots,t_{N}$, where $T$ is represented using a negative Schur form network, $B$ is represented using a low-rank bilinear form network, and $C$ and $g$ are represented using feed-forward neural networks. The trajectories are then transformed back into the original coordinate system using
$z_{n}=h^{-1}\left(\left[\begin{array}[]{cc}y_{n}&x_{n}\end{array}\right]^{\rm T};\epsilon\right),\qquad n=1,\dots,N.$ (15)
The key components of Definition 3, including the invertible coupling flow network, Schur form network, and low-rank bilinear form network, will be defined in the following sections. First we will begin by stating the main result of the paper.
@@
Due to Theorem 2, normally stable fast-slow systems can be represented using (7)–(2). Therefore, the FSNN inherits the universal approximation properties of its components, which we will later show are universal approximators for their corresponding class of functions. The invertible coupling flow network which represents $h$ is described in Section 3.1 and was proven to be a universal approximator to invertible maps in [58]. The universal approximation theorems of the negative Schur form network representing $T$ and the bilinear form network representing $B$ are proved in Sections 3.2 and 3.3, respectively. The universal approximation of nonlinear functions using feedforward neural networks for $C$ and $g$ are well known, e.g. [15]. Finally, we describe a consistent numerical scheme to obtain trajectories in time, which has an error bound that vanishes as the time steps $\Delta t_{n}=t_{n}-t_{n-1}\rightarrow 0$.

### v2

@@ Abstract
Singularly perturbed dynamical systems play a crucial role in climate dynamics and plasma physics. A powerful and well-known tool to address these systems is the Fenichel normal form, which significantly simplifies fast dynamics near slow manifolds through a transformation. However, this normal form is difficult to realize in conventional numerical algorithms. In this work, we explore an alternative way of realizing it through structure-preserving machine learning. Specifically, a fast-slow neural network (FSNN) is proposed for learning data-driven models of singularly perturbed dynamical systems with dissipative fast timescale dynamics. Our method enforces the existence of a trainable, attracting invariant slow manifold as a hard constraint. Closed-form representation of the slow manifold enables efficient integration on the slow time scale and significantly improves prediction accuracy beyond the training data. We demonstrate the FSNN on examples including the Grad moment system, two-scale Lorenz96 equations, and Abraham-Lorentz dynamics modeling radiation reaction of electrons.
@@
Singularly perturbed dynamical systems, slow manifold, model reduction, closures
@@ 1 Introduction
In this article, a new architecture called the Fast-Slow Neural Network (FSNN) is developed for learning data-driven models of singularly perturbed dynamical systems with dissipative fast timescale dynamics. We refer to our approach as an intelligent attractor since the method enforces the existence of a trainable, attracting invariant slow manifold [1] as a hard constraint. The constraint is imposed by way of a novel neural network architecture with structured weight matrices. By leveraging the neural ODE framework from [2], our technique produces deterministic evolution equations that approximately reproduce trajectory training data. When the training data comes from an unknown dynamical system, the method achieves data-driven full-order model discovery with stability guarantees due to provable existence of the slow manifold. In short, our architecture is, by design, a singularly perturbed dynamical system which contains an attractor.
@@
Nonlinear dimensionality reduction inspired by inertial manifold theory has a rich history (see [5] and the references therein) originating in the 1980s and 1990s [6, 4, 7, 8, 9, 10, 11, 12]. Data-driven approaches for discovering inertial manifolds in dynamical systems and using separations of scales to evolve a lower-dimensional reduced order model first materialized in the 90s [13, 14, 15, 16, 17, 18, 19] and recently reemerged in popularity due to boom in machine-learning-based modeling [20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32]. The work in this manuscript tackles a special class of systems with attracting invariant sets, namely fast-slow systems with dissipative fast timescale dynamics [1]. Unlike previous data-driven approaches, the FSNN is a full-order model which, using neural network architectures with special properties, strictly enforces the existence of an attracting slow manifold. This method may be viewed as a major step along a pathway in developing data-driven model discovery techniques with embedded invariant manifolds. In future work, we plan to address the general case where explicit timescale separation is not used as a crutch.
@@
The design of the FSNN architecture is motivated by the first few theorems discovered by Fenichel [36]. The components of the architecture are constructed using structured weight matrices. The coordinate transformation into Fenichel normal form is based on an invertible coupling flow network [37] in which we introduced the additional structure-preserving property of bi-Lipschitz control through the use of bi-Lipschitz affine transformation (bLAT) layers. This network and its inverse can be evaluated using closed-form expressions with identical computational complexity. We note that “vanilla” neural networks are generally non-invertible [38, 39, 40]. We also introduce the Schur form network, which is a novel architecture based on the Schur decomposition for controlling the locations of eigenvalues in the complex plane, to parameterize the linear term in the dynamics. Additionally, a low-rank bilinear map network is included to increase the approximation power in the fast variable space. Our approach is then integrated using adaptive time stepping that treats the linear term implicitly, further leveraging back-substitution on the Schur form. Some of the structures introduced are related to the idea of trivializing manifold optimization through parameterizations [41]. Structure-preserving flow maps through machine learning are an active area of research, and include such as the symplectic neural networks for Hamiltonian dynamics [42, 43]. Previous studies on flow map learning for multi-scale dynamical systems include [44] and [45]. In particular, [45] simultaneously preserves symplecticity and adiabatic invariance in a multi-scale setting.
Before going into a detailed presentation of our new method, it is worth highlighting an important shortcoming it suffers from, and how that shortcoming might be mitigated in practice. The method assumes that the dimension of the attracting invariant manifold is known in advance. While this is not a limitation in the surrogate modeling context, where the (expensive to simulate) dynamical system is known in detail, it does present a moderate challenge in the model discovery context. As such, when performing data-driven model discovery using our method, it will be necessary to first determine the dimension of the slow manifold, either through physical reasoning, or through application of auxiliary data mining techniques such as diffusion maps [46], autoencoders [26], or other techniques from reduced order modeling [5]. For a review on various manifold learning techniques, we refer the reader to [47]. In many cases, including the examples considered in this manuscript, the form of the slow manifold and therefore its dimensionality are known in the singular limit $\epsilon=0$ as a power series expansion of $\epsilon$. However, the exact form for $\epsilon>0$ and the transformation into Fenichel normal form is unknown. In order to learn the slow manifold effectively, a sufficient number of trajectories for various choices of $\epsilon$ and initial conditions are necessary. In the final example of the paper, we apply our network in the nontrivial setting of learning the slow manifold for the Abraham-Lorentz equations which, for trajectories starting off of the slow manifold, are unstable forward in time.
@@ 2 Fast-Slow Dynamics and the Fenichel Normal Form
In this paper, we develop a novel machine learning (ML)-inspired architecture, the Fast-Slow Neural Network (FSNN). FSNNs are a parameterization for a class of dynamical systems exhibiting multiscale behavior in time that, to leading order, are dissipative on the fast time scale. Fast-slow systems, which exhibit both short-term and long-term behavior, are relevant to many physical systems observed in nature, including the climate and plasma systems. In this section, we motivate the FSNN architecture with a review of the theory of fast-slow systems.
@@
A fast-slow system is an ODE of the form [36, 1]
(1a) $\epsilon\frac{d}{d\tau}y$ $=f(x,y,\epsilon),$
(1b) $\frac{d}{d\tau}x$ $=g(x,y,\epsilon),$
@@
(2) $\epsilon\frac{d}{d\tau}z$ $=U(z,\epsilon),$
@@
Here a bounded hypersurface is called a locally invariant manifold if the vector field defining the dynamical system is tangent to the hypersurface at all of its points [48]. This lemma is proved following Theorem 4 in [36]. The purpose of the FSNN is to be a universal approximator for normally stable fast-slow systems. The network is based on the following well known result due to Fenichel.
@@
(4a) $\frac{d}{dt}\tilde{y}$ $=A(\tilde{x},\tilde{y},\epsilon)\tilde{y},$
(4b) $\frac{d}{dt}\tilde{x}$ $=\epsilon{\tilde{g}}(\tilde{x},\tilde{y},\epsilon),$
@@
The above lemma is proved in Chapter 3.2 of [36]. The lemma introduces (4), which is often called the Fenichel normal form due to the pioneering work of Neil Fenichel [3]. Note that the Fenichel normal form given in [36] is more general than (4), including fixed points that are either stable or unstable in the linearization of the fast variable. In this work, we only consider fast variables that are long-time stable. Moreover, due to the existence of the coordinate transformation, we consider the general form of the singularly perturbed systems (2). However, the advantage of the split form (1) is that the slow manifold dimension is known.
@@
(7) $\left[\begin{array}[]{c}y\\ x\end{array}\right]$ $=h(z;\epsilon),$
@@
(8a) $\frac{d}{dt}y$ $=T(x)y+B(x,y,\epsilon)(y,y)+\epsilon C(x,y,\epsilon)y,$
(8b) $\frac{d}{dt}x$ $=\epsilon g(x,y,\epsilon).$
$T(x)\in\mathbb{R}^{N_{y}\times N_{y}}$ is a block upper triangular matrix consisting of either $\mathbb{R}^{2\times 2}$ or $\mathbb{R}$ blocks on the diagonal and the eigenvalues of these diagonal blocks are strictly to the left of the imaginary axis for all $x$. $B(x,y,\epsilon):\mathbb{R}^{N_{y}}\times\mathbb{R}^{N_{y}}\rightarrow\mathbb{R}^{N}_{y}$ is a bilinear map, $C(x,y,\epsilon)\in\mathbb{R}^{N_{y}\times N_{y}}$ is a matrix, and $g(x,y,\epsilon)\in\mathbb{R}^{N_{x}}$. $T$, $B$, $C,$ and $g$ are $C^{r}$ for $r\geq 1$.
In the following discussion, we will prove Theorem 8 by refining the Fenichel normal form in (4) and applying the matrix Schur decomposition. First, we modify (4) by performing a Taylor expansion of the term $A(\tilde{x},\tilde{y},\epsilon)\tilde{y}$ about $\tilde{y}=0$ and $\epsilon=0$. Assuming that $A$ is in $C^{r}$ for $r\geq 1$, we have
(9) $A(\tilde{x},\tilde{y},\epsilon)\tilde{y}=A(\tilde{x},0,0)\tilde{y}+B(\tilde{x},\tilde{y},\epsilon)(\tilde{y},\tilde{y})+\epsilon C(\tilde{x},\tilde{y},\epsilon)\tilde{y},$
@@
The Schur form of a matrix, $A\in\mathbb{R}^{M\times M}$, is a matrix $T\in\mathbb{R}^{M\times M}$ that forms the Schur decomposition $A=QTQ^{T}$, where $Q\in\mathbb{R}^{M\times M}$ is orthogonal ($Q^{T}Q=I$) and $T$ is block upper triangular where the diagonal blocks consist of $\mathbb{R}$ and $\mathbb{R}^{2\times 2}$ blocks.
The existence of the Schur decomposition is proven in many references, including [49, 50], though for any given matrix, the decomposition is not unique. However, the eigenvalues of $A$ are given by the eigenvalues of the diagonal blocks of $T$. Let the Schur decomposition of $A(\tilde{x},0,0)$ be given by
(10) $A(\tilde{x},0,0)=Q(\tilde{x})T(\tilde{x})Q(\tilde{x})^{\rm T},$
@@
where $\tilde{B}(\tilde{x},\hat{y},\epsilon)$ is a bilinear map and $\tilde{C}(\tilde{x},\hat{y},\epsilon)$ is a linear map. This establishes an equivalence to (8).
@@
(11) $\frac{d}{d\tau}x=g(x,0,\epsilon),\qquad(x,y)\in M,$
@@
(12) $d\left((x(t),y(t)),M\right)\leq\kappa\exp(\alpha t),$
@@ 3 Fast-Slow Neural Networks (FSNNs)
The fast slow neural network (FSNN) is a transformation of an initial condition, $z_{0}$, and parameter, $\epsilon>0$, to a time series $\{z_{n}\}_{n=0}^{N}$ corresponding to times $t_{0},\dots,t_{N}$, where $z_{n}\in\mathbb{R}^{N}$ for $n=0,\dots,N$. The initial condition is transformed into fast-slow coordinates, $(y_{0},x_{0})$, via (7) where $h$ is represented using an invertible coupling flow network. A numerical integration is used to approximate the solution to the fast-slow system in (8) to obtain the time series $\{(y_{n},x_{n})\}_{n=0}^{N_{t}}$ for corresponding times $t_{0},\dots,t_{N}$, where $T$ is represented using a negative Schur form network, $B$ is represented using a low-rank bilinear map network, and $C$ and $g$ are represented using feed-forward neural networks. The trajectories are then transformed back into the original coordinate system using
(14) $z_{n}=h^{-1}\left(\left[\begin{array}[]{cc}y_{n}&x_{n}\end{array}\right]^{\rm T};\epsilon\right),\qquad n=1,\dots,N.$
The key components of Definition 3, including the invertible coupling flow network, Schur form network, and low-rank bilinear map network, will be defined in the following sections. First we will begin by stating the main result of the paper.
@@
Due to Theorem 8, normally stable fast-slow systems can be represented using (7)–(8). Therefore, the FSNN inherits the universal approximation properties of its components, which we will later show are universal approximators for their corresponding class of functions. The invertible coupling flow network which represents $h$ is described in Section 3.1 and was proven to be a universal approximator to invertible maps in [37]. The universal approximation theorems of the negative Schur form network representing $T$ and the bilinear map network representing $B$ are proved in Sections 3.2 and 3.3, respectively. The universal approximation of nonlinear functions using feedforward neural networks for $C$ and $g$ are well known, e.g. [52]. Finally, we describe a consistent numerical scheme to obtain trajectories in time, which has an error bound that vanishes as the time steps $\Delta t_{n}=t_{n}-t_{n-1}\rightarrow 0$.